1. Rate the request on scientific merit. 5 This research is original and is important to the scientific community. The researchers have a firm understanding of previous work done in the field. They seem to do excellent research, and their number of publications is extremely good.

2. Rate the request on potential for significant progress. 4 I believe the researchers have excellent experience in this field and should be able to meet most of their goals. They have a very large set of goals, and might have trouble reaching all of their goals in one and progress has been made from the previous year. Since they have already published many results from the previous years, I believe that they should have no trouble keeping up with publishing good results.

3. Rate the request on appropriateness of SDSC resources. 2 This project will determine new and interesting results for the research being done. My problem with their proposal is that they do not even mention autotasking, nor do they mention about running extremely large scale runs. It is my belief that a good workstation of today could easily satisfy their needs. I think that we need to make a clear indication that supercomputing is computing large scale runs with multiple processors. They mention results of running at 100 Mflops on the C90 for some of their runs, and they mention the words “optimized”. I am puzzled since I have worked with many codes, and on a C90, on a single CPU, I have been able to achieve in the 500Mflop range for non-trivial codes. I realize that their runs take many hours on a C90, which clearly indicates that parallelization of their code is necessary.

4. How many Service Units do you recommend be allocated? 450 This is an extremely good proposal for some purposes, but not for supercomputing! Using 25 hours of C90 time per run, without talking about autotasking is a clear indication that these researchers are doing supercomputing in the late 80’s. This is the mid 90’s, where PC’s are now using multiple CPU’s. Research in supercomputing should help advance computing as well as the research in their field. Another problem that I have with their research is that they are requesting 1000 SU’s, and yet they require 4000SU’s. If they can accomplish most of their research without a C90, then clearly they can accomplish all of their research without this machine. They do not mention which runs they will try to accomplish on this machine. I recommend that they get funded for their 2D runs, (200SU) and their 1D runs (250SU). I think that their 3D runs should be implemented on their workstations, (they could use the machine just for these runs and probably get a faster turnaround on these machines compared to waiting in the batch for these runs to finish.)